OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2.

OpenStudy (lgbasallote):

the slope-intercept form is y = mx + b where: m = slope b = y-intercept so just substitute your slope and y-intercept into y = mx + b does that help?

OpenStudy (anonymous):

kind of i still dont understand how i would write it out

OpenStudy (lgbasallote):

let me give an example slope is 2 and y-intercept is -1 i would write it as y = mx + b in my example slope is 2 so m = 2 and y-intercept is -1 so b= -1 so i sub y = 2x + -1 or simply y= 2x - 1 does that help?

OpenStudy (saifoo.khan):

@lgbasallote http://i1.kym-cdn.com/entries/icons/original/000/007/423/untitle.JPG